# Stability theory over toroidal or Novikov-type base and canonical modifications

### v2

@@ 2.2. Generalized test configurations and examples
  2. (2)
    For an affine integral reduced scheme $Y$ of finite type over a field $\mathbbm{k}$ (resp., polarized variety $(Y,L)$ over $\mathbbm{k}$), an (affine) algebraic test deformation (resp., (polarized) algebraic test deformation of $(T\curvearrowright Y,\xi)$), where $T$ is an algebraic $\mathbbm{k}$-torus $N\otimes\mathbb{G}_{m}$ and $\xi$ is an element of $N\otimes\mathbb{R}$, is a generalized (unpolarized affine) test configuration $p\colon\mathcal{Y}\to U_{\tau}$ with $p^{-1}(0)=Y$.
    Note that if $\mathbbm{k}$ contains $\mathbb{R}$ and $T$ is $\mathbbm{k}$-split, $\xi$ can be identified with its holomorphic333we take this choice rather than real vector field, as in usual choice of Reeb vector field, as we want to work also over $\mathbbm{k}\neq\mathbb{C}$ vector field $\xi$ over the smooth locus $Y^{\rm sm}$ of $Y$. Indeed, if we consider the induced action of $T$ on $p^{-1}(0)$ and the natural map $a\colon N_{\mathbb{R}}\to H^{0}(p^{-1}(0),T_{p^{-1}(0)^{\rm sm}})$ induced by it (here, $T_{p^{-1}(0)^{\rm sm}}$ denotes the Zariski tangent sheaf of the smooth locus of $p^{-1}(0)$), the image ${\rm Im}(a)$ contains $a(\xi)$ and $a$ is often injective, so that we identify $\xi$ with a holomorphic vector field.
  3. (3)
@@
For any $\xi=(w_{1},\cdots,w_{r})\in\mathbb{R}^{r}$, we take a lattice $N$ which is the minimum sublattice of $\mathbb{Z}^{r}$ such that $N\otimes\mathbb{R}(\subset\mathbb{R}^{r})$ which contains $\mathbb{R}\xi$. Then the natural pairing map
@@
We apply the above for our weight vector $\xi=(w_{1},\cdots,w_{r})\in\mathbb{R}^{r}$, and also take a rational polyhedral cone $\tau$ of $N$ which contains $\xi$, to later discuss its perturbations inside $\tau$.
@@
We end Example 2.9 by the following conclusion of our discussion.
@@
For any affine klt variety $X$ with its closed point $x$, over an algebraically closed field $\mathbbm{k}$ of characteristic 00, consider the volume-minimizing valuation $v_{X}$ of rational rank $r$ ([Li18, Blu18, Xu20]) and $W:=\mathop{\mathrm{Spec}}\nolimits({\rm gr}_{v_{X}}(\mathcal{O}_{X,x}))$. Take the groupificiation $N$ of ${\rm Im}(v_{X})$ and its small enough rational polyhedral cone $\tau$ which includes $\xi=(w_{1},\cdots,w_{l})$ where $w_{i}$ are the values of $v_{X}$ for homogeneous generators of ${\rm gr}_{v_{X}}(\mathcal{O}_{X,x})$.
@@
Recall that $C$ is given an action of algebraic torus $T$ from the definition, and by the reductivity of $G:={\rm Aut}_{T}(C)$ ([DS17, Appendix], cf., also [Od24a]), which means the group of automorphisms of $C$ which commutes with the $T$-action, and the action of its centralizer to the multi-Hilbert scheme [HM04] $H$ as in [DS17]. From these facts, it is easy to see similar remark apply as above Remark 2.12.
@@
  5. (5)
    (as a convention: left-continuous) $\cap_{\lambda^{\prime}\leq\lambda}\mathcal{F}^{\lambda^{\prime}}R_{m}=\mathcal{F}^{\lambda}R_{m}$.
@@
Conlon-Hein [CH22] introduces a notion of negative weight deformation of Calabi-Yau cones in terms of sequence of affine test configurations, for their classification of asymptotically conical Calabi-Yau manifolds. We alternatively introduce the notion of algebraic negative deformation in Definition 2.6 without using sequence (to answer Open Questions 1.4(2) of [CH22]). For our later purpose, we also clarify the relation as follows. Here, we assume $\mathbbm{k}=\mathbb{C}$ as [CH22]. We take a affine $\mathbb{C}$-variety $Y$ and an algebraic torus $T=N\otimes\mathbb{G}_{m}$ acting on $Y$, with $\xi\in N_{\mathbb{R}}$.
@@
For the item (1), we can assume that $\tau=\mathbb{Z}_{\geq 0}^{r}$, $N=\mathbb{Z}^{r}$. Then we approximate $\xi$ by $\xi_{i}=(\frac{r_{i,1}}{m_{1}},\frac{r_{i,2}}{m_{1}},\cdots,\frac{r_{i,r}}{m_{1}})\in\tau\cap N_{\mathbb{Q}}$, with nonnegative integers $r_{i,j}(j=1,\cdots,r)$ and $m_{i}$s so that a convergence $\xi_{i}\to\xi$ holds for $i\to+\infty$. Then take a morphism $\varphi_{\vec{r}}\colon\mathbb{A}^{1}\to U_{\tau}$ defined by $t\mapsto(t^{r_{i,1}},t^{r_{i,2}},\cdots,t^{r_{i,r}})$ and consider the base change $\mathcal{Y}\times_{p,U_{\tau},\varphi_{\vec{r}}}\mathbb{A}^{1}$, which we denote as $p^{\prime}_{i}\colon W^{\prime}_{i}\to\mathbb{A}_{1}$. It can be naturally seen as an affine test configuration but we modify the $\mathbb{G}_{m}$-action $\sigma^{\prime}_{i}$ on it by compositing with the $m_{i}$-th power map $\mathbb{G}_{m}\to\mathbb{G}_{m}$. In particular, $p^{\prime}_{i}(\sigma^{\prime}_{i}(z,x))=z^{m_{i}}p^{\prime}_{i}(x)$ for $z\in\mathbb{G}_{m}(\mathbbm{k}),x\in W^{\prime}_{i}$. Then, we take the base change of $p^{\prime}_{i}\colon W^{\prime}_{i}\to\mathbb{A}^{1}$ with $\sigma^{\prime}_{i}$ by the $m_{i}$-th power map $q_{i}\colon\mathbb{A}^{1}\to\mathbb{A}^{1}$ whose target is regarded as the base of $p^{\prime}_{i}$. It is straightforward confirmation that the sequence, with respect to $i=1,2,\cdots$, of the obtained fiber product $p_{i}\colon W_{i}(:=W^{\prime}_{i}\times_{p^{\prime}_{i},\mathbb{A}^{1},q_{i}}\mathbb{A}^{1})\to\mathbb{A}^{1}$ with the $\mathbb{G}_{m}$-action induced by $\sigma^{\prime}_{i}$ satisfies the conditions of the [CH22, Definition 1.7], with $c_{i}$ of loc.cit as $m_{i}$, $\mu_{i}$ of loc.cit as $1$, hence the latter being auxiliary parameters. Suppose $Z$ is the maximum Artin scheme supported at the $T$-invariant point $p_{\tau}\in U_{\tau}$, such that $\mathcal{Y}|_{Z}$ is trivial family. Such $Z$ exists due to the Zorn’s lemma combined with the formal smoothness ([Gro64, Part I, Part IV]) of ${\rm Aut}_{T}(Y)$ which is an algebraic group as the stabilizer of $[Y]$ in the multi-Hilbert scheme. By using this $Z$, $k_{i}$ of loc.cit can be written as $m_{i}{\rm length}\mathcal{O}_{\varphi_{\vec{r}}^{*}Z}$, which explains the nature of the $\xi$-weight in loc.cit. ∎
The item (2) is proved in [Od24c] in which some more related materials are also explored.
@@ 3.1. Toroidal stack-theoretic setup
    generalizing the case $r=1,\tau=\mathbb{R}_{\geq 0}$ treated in [HL14, AHLH23], which is denoted simply as $\Theta$. We denote the $T$-invariant point of $U_{\tau}$ as $p_{\tau}$. In [HL14], (the isomorphic object) $\Theta_{\tau}$ for general $r$ and regular simplicial $\tau$ is also introduced under the disguise of $\Theta^{r}$, the $r$-times self-product of $\Theta$. For our purpose, since we will encounter a setup with various $\tau$ for a fixed $N$ later, we prefer the above setup and notation.
  2. (2)
    We fix a valuation ring $(R,v)$ of height $1$ with $\Gamma_{\geq 0}:=v(R\setminus\{0\})\subset\mathbb{R}$. Further, take a polyhedral cone $\tau^{\prime}$ of $N\oplus\Gamma_{\geq 0}$ which projects as PL isomorphism (in particular, bijectively) to $\tau\subset N_{\mathbb{R}}$ and the normal vector of $\tau$ can be taken as $(\vec{m},1)\in M\oplus(\Gamma_{\geq 0})^{*}$, i.e., $\langle\tau,(\vec{m},1)\rangle=0$. We denote the character on $T$ which corresponds to $\vec{m}$ as $\chi_{\vec{m}}$. Then we define
@@
again generalizing the case $r=1,\tau=\mathbb{R}_{\geq 0}$ introduced in [Hei18] (cf., also [AHLH23]). Note that the adjunction of $\frac{\pi}{\chi_{\vec{m}}}$ in the right hand side of (1) is a localization corresponding to a “toroidal” blow up of $R\times U_{\tau}$, with respect to $\hat{\tau}^{\prime}:=\tau^{\prime}+\mathbb{R}_{\geq 0}(\vec{0},1)$. In particular, we have a natural (birational) map $b\colon{\overline{\rm ST}}_{R}(\tau^{\prime})\to{\rm Spec}(R)$ and a canonical open embedding ${\rm Spec}(K)\hookrightarrow{\overline{\rm ST}}_{R}(\tau^{\prime}),$ where $K$ denotes the fraction field of the valuation ring $R$.
@@
Definition 3.3 (Higher $\Theta$-strata cf., [HL14, §2.1]).
Consider the mapping stack ${\rm Map}(\Theta^{r},\mathcal{M})$ of $\Theta^{r}$ to $\mathcal{M}$, as an algebraic stack, after [AHR20, 5.10, 5.11], [AHR19, 6.23], [HLP23, 5.1]. Extending [HL14, Definition 2.1.1], a higher $\Theta$-stratum of rank $r$ in $\mathcal{M}$ consists of a union of connected components $\mathcal{Z}^{+}\subset{\rm Map}(\Theta^{r},\mathcal{M})$ such that the natural evaluation morphism ${\rm ev}_{(1,\cdots,1)}\colon\mathcal{Z}^{+}\to\mathcal{M}$ is a closed embedding. If we fix an isomorphism $\Theta^{r}\simeq\Theta_{\tau}$ for a rational polyhedral cone $\tau$, then we also call higher $\Theta$-stratum as higher $\Theta$-stratum for the cone $\tau$.
Further, if $\mathcal{M}$ is quotient stack of the form $[M/G]$ where $M$ is an algebraic $\mathbbm{k}$-scheme $M$ and $G$ is an algebraic $\mathbbm{k}$-group, we often write ${\rm ev}_{(1,\cdots,1)}(\mathcal{Z}^{+})=[Z^{+}/G]$ for some $G$-invariant locally closed subscheme $Z^{+}\subset M$.
Definition 3.4 (Elementary ($\tau$-)modification cf., [AHLH23, §3.5, Definition 3.36]).
An elementary ($\tau$-)modification of $f\colon{\rm Spec}(R)\to\mathcal{M}$ is the data of a morphism $\overline{f}\colon\overline{\rm ST}_{R}(\tau)\to\mathcal{M}$ along with an isomorphism $f\simeq\overline{f}|_{\rm Spec(R)}$. Here, to make sense of the restriction symbol, $\rm Spec(R)$ is regarded as a localization of $\overline{{\rm ST}}_{R}(\tau^{\prime})$ since we have $\mathop{\mathrm{Spec}}\nolimits(R)=[\mathop{\mathrm{Spec}}\nolimits R[M]/T]$.
@@
Suppose $(R,v=v_{R})$ is a valuation local ring, such that the associated graded ring $gr_{v}(R):=\oplus_{w\in{\rm Im}(v)}\{a\in R\mid v(a)\geq w\}/\{a\in R\mid v(a)>w\}$ is isomorphic to a semigroup ring over a field $\mathbbm{k}$ and a submonoid of $\mathbb{R}^{r}$ with some $r\in\mathbb{Z}_{>0}$ as a graded ring (which holds e.g. if $R$ is of the form $\mathbbm{k}[[\Gamma_{\geq 0}]]$ for a field $\mathbbm{k}$ in Definition 2.1) and $\xi,N,M_{\xi\geq 0}$ are as Notation 1. Now, we set
$R[[M_{\xi\geq 0}]]:=\{\sum_{i=1,2,\cdots}a_{i}t^{m_{i}}\mid a_{i}\in R,m_{i}\in M_{\xi\geq 0},m_{1}<m_{2}<\cdots,m_{i}\to\infty(i\to\infty)\},$
@@
3.2. Higher $\Theta$-stable reduction theorem
Below, we generalize the semistable reduction type theorem in [AHLH23, §6.1], which was for $r=1$ case. Morally speaking, it says that a one parameter family which strictly degenerates into a higher $\Theta$-strata, there is a canonical modification of the family to avoid a higher $\Theta$-strata i.e., with “less degeneration”. This is a version of stable reduction type theorem, and later Theorem 3.18 is its iterated version along a certain kind of stratification (higher $\Theta$-stratification cf., Definition 3.17).

### v3

@@ 2.2. Generalized test configurations and examples
  2. ($\mathrm{ii}$)
    For an affine integral reduced scheme $Y$ of finite type over a field $\mathbbm{k}$ (resp., polarized variety $(Y,L)$ over $\mathbbm{k}$), an (affine) algebraic test deformation (resp., (polarized) algebraic test deformation of $(T\curvearrowright Y,\xi)$), where $T$ is an algebraic $\mathbbm{k}$-torus $N\otimes\mathbb{G}_{m}$ i.e., which has the co-character lattice $N$, and $\xi$ is an element of $N\otimes\mathbb{R}$, is a generalized (unpolarized affine) test configuration $p\colon\mathcal{Y}\to U_{\tau}$ with $p^{-1}(0)=Y$.
    Note that if $\mathbbm{k}$ contains $\mathbb{R}$ and $T$ is $\mathbbm{k}$-split, $\xi$ can be identified with its holomorphic444we take this choice rather than real vector field, as in usual choice of Reeb vector field, as we want to work also over $\mathbbm{k}\neq\mathbb{C}$ vector field $\xi$ over the smooth locus $Y^{\rm sm}$ of $Y$. Indeed, if we consider the induced action of $T$ on $p^{-1}(0)$ and the natural map $a\colon N_{\mathbb{R}}\to H^{0}(p^{-1}(0),T_{p^{-1}(0)^{\rm sm}})$ induced by it (here, $T_{p^{-1}(0)^{\rm sm}}$ denotes the Zariski tangent sheaf of the smooth locus of $p^{-1}(0)$), the image ${\rm Im}(a)$ contains $a(\xi)$ and $a$ is often injective, so that we identify $\xi$ with a holomorphic vector field.
  3. ($\mathrm{iii}$)
@@
For any $\xi=(w_{1},\cdots,w_{l})\in\mathbb{R}^{l}$, we take a lattice $N$ which is the minimum sublattice of $\mathbb{Z}^{l}$ such that $N\otimes\mathbb{R}(\subset\mathbb{R}^{l})$ which contains $\mathbb{R}\xi$. Then the natural pairing map
@@
We apply the above for our weight vector $\xi=(w_{1},\cdots,w_{l})\in\mathbb{R}^{l}$, and also take a rational polyhedral cone $\tau$ of $N$ which contains $\xi$, to later discuss its perturbations inside $\tau$.
@@
We end Example 2.9 by the following brief summary of our discussion.
@@
For any affine klt variety $X$ with its closed point $x$, over an algebraically closed field $\mathbbm{k}$ of characteristic 00, consider the volume-minimizing valuation $v_{X}$ of $\mathbb{Q}$-rank $r$ in the sense of [Li18], and $W:=\mathop{\mathrm{Spec}}\nolimits({\rm gr}_{v_{X}}(\mathcal{O}_{X,x}))$.
There is an affine toric variety $U_{\tau}$ and an algebraic positive weight deformation over it
$\pi_{\tau}\colon\mathcal{X}_{\tau}\twoheadrightarrow U_{\tau},$
which restricts to $W\times\partial U_{\tau}\twoheadrightarrow\partial U_{\tau}$, over the toric boundary $\partial U_{\tau}:=U_{\tau}\setminus T$. Its general fibers are $X$.
@@
Recall that $C$ is given an action of algebraic split torus $T$ from the definition, and by the reductivity of $G:={\rm Aut}_{T}(C)$ ([DS17, Appendix], cf., also [Od24a]), which means the group of automorphisms of $C$ which commutes with the $T$-action, and the action of its centralizer to the multi-Hilbert scheme [HM04] $H$ as in [DS17]. From these facts, it is easy to see similar remark apply as above Remark 2.12.
@@
  5. ($\mathrm{v}$)
    (as a convention: left-continuous) $\cap_{\lambda^{\prime}<\lambda}\mathcal{F}^{\lambda^{\prime}}R_{m}=\mathcal{F}^{\lambda}R_{m}$.
@@
Conlon-Hein [CH22] introduces a notion of negative weight deformation of Calabi-Yau cones in terms of sequence of affine test configurations, for their classification of asymptotically conical Calabi-Yau manifolds. We alternatively introduce the notion of algebraic negative deformation in Definition 2.6 without using sequence (to answer Open Questions 1.4(2) of [CH22]). For our later purpose, we also clarify the relation as follows. Here, we assume $\mathbbm{k}=\mathbb{C}$ as [CH22]. We take a affine $\mathbb{C}$-variety $Y$ and an algebraic split torus $T=N\otimes\mathbb{G}_{m}$ acting on $Y$, with $\xi\in N_{\mathbb{R}}$.
@@
For the item ($\mathrm{i}$), we can assume that $\tau=\mathbb{Z}_{\geq 0}^{r}$, $N=\mathbb{Z}^{r}$. Then we approximate $\xi$ by $\xi_{i}=(\frac{r_{i,1}}{m_{1}},\frac{r_{i,2}}{m_{1}},\cdots,\frac{r_{i,r}}{m_{1}})\in\tau\cap N_{\mathbb{Q}}$, with nonnegative integers $r_{i,j}(j=1,\cdots,r)$ and $m_{i}$s so that a convergence $\xi_{i}\to\xi$ holds for $i\to+\infty$. Then take a morphism $\varphi_{\vec{r}}\colon\mathbb{A}^{1}\to U_{\tau}$ defined by $t\mapsto(t^{r_{i,1}},t^{r_{i,2}},\cdots,t^{r_{i,r}})$ and consider the base change $\mathcal{Y}\times_{p,U_{\tau},\varphi_{\vec{r}}}\mathbb{A}^{1}$, which we denote as $p^{\prime}_{i}\colon W^{\prime}_{i}\to\mathbb{A}_{1}$. It can be naturally seen as an affine test configuration but we modify the $\mathbb{G}_{m}$-action $\sigma^{\prime}_{i}$ on it by compositing with the $m_{i}$-th power map $\mathbb{G}_{m}\to\mathbb{G}_{m}$. In particular, $p^{\prime}_{i}(\sigma^{\prime}_{i}(z,x))=z^{m_{i}}p^{\prime}_{i}(x)$ for $z\in\mathbb{G}_{m}(\mathbbm{k}),x\in W^{\prime}_{i}$. Then, we take the base change of $p^{\prime}_{i}\colon W^{\prime}_{i}\to\mathbb{A}^{1}$ with $\sigma^{\prime}_{i}$ by the $m_{i}$-th power map $q_{i}\colon\mathbb{A}^{1}\to\mathbb{A}^{1}$ whose target is regarded as the base of $p^{\prime}_{i}$. It is straightforward confirmation that the sequence, with respect to $i=1,2,\cdots$, of the obtained fiber product $p_{i}\colon W_{i}(:=W^{\prime}_{i}\times_{p^{\prime}_{i},\mathbb{A}^{1},q_{i}}\mathbb{A}^{1})\to\mathbb{A}^{1}$ with the $\mathbb{G}_{m}$-action induced by $\sigma^{\prime}_{i}$ satisfies the conditions of the [CH22, Definition 1.7], with $c_{i}$ of loc.cit as $m_{i}$, $\mu_{i}$ of loc.cit as $1$, hence the latter being auxiliary parameters. Suppose $Z$ is the maximum Artin scheme supported at the $T$-invariant point $p_{\tau}\in U_{\tau}$, such that $\mathcal{Y}|_{Z}$ is trivial family. Such $Z$ exists due to the Zorn’s lemma combined with the formal smoothness ([Gro64, Part I, Part IV]) of ${\rm Aut}_{T}(Y)$ which is an algebraic group as the stabilizer of $[Y]$ in the multi-Hilbert scheme. By using this $Z$, $k_{i}$ of loc.cit can be written as $m_{i}{\rm length}\mathcal{O}_{\varphi_{\vec{r}}^{*}Z}$, which explains the nature of the $\xi$-weight in loc.cit. ∎
The item ($\mathrm{ii}$) is proved in [Od24c] in which some more related materials are also explored.
@@ 3.1. Toroidal stack-theoretic setup
    generalizing the case $r=1,\tau=\mathbb{R}_{\geq 0}$ treated in [HL14, AHLH23], which is denoted simply as $\Theta$. We denote the $T$-invariant point of $U_{\tau}$ as $p_{\tau}$. In [HL14], (the isomorphic object) $\Theta_{\tau}$ for the case of regular simplicial $\tau$ is also essentially introduced under the disguise of $\Theta^{r}$, the $r$-times self-product of $\Theta$. For our purpose, since we will encounter a setup with various $\tau$ for a fixed $N$ later, we prefer the above setup and notation.
  2. ($\mathrm{ii}$)
    We fix a valuation ring $(R,v)$ of height $1$ with $\Gamma_{\geq 0}:=v(R\setminus\{0\})\subset\mathbb{R}$. To set up a notion of elementary modification of family along $R$ (Definition 3.4 later), we prepare the following.
    Further, we set $N_{\mathbb{R}}:=N\otimes_{\mathbb{Z}}\mathbb{R}$ and take a $r$-dimensional polyhedral cone $\tau^{\prime}$ inside $N_{\mathbb{R}}\oplus\mathbb{R}$, whose generators can be taken inside $N\times\Gamma_{\geq 0}$ and assume that the restriction of the first projection $N_{\mathbb{R}}\times\mathbb{R}\to N_{\mathbb{R}}$ to $\tau^{\prime}$ is a PL isomorphism (in particular, bijectively) to $\tau\subset N_{\mathbb{R}}$. In other words, $\tau^{\prime}$ is the graph of a homomorphism $\tau\to\mathbb{R}_{\geq 0}$.
    We further assume the normal vector of $\tau$ can be taken as $(\vec{m},1)\in M\oplus(\Gamma_{\geq 0})^{*}$, i.e., $\langle\tau,(\vec{m},1)\rangle=0$. We denote the character on $T$ which corresponds to $\vec{m}$ as $\chi_{\vec{m}}$. Then we define
@@
again generalizing the case $r=1,\tau=\mathbb{R}_{\geq 0}$ introduced in [Hei18, §2.B] and [AHLH23, §3.5]. Note that, if $R=\mathbbm{k}[[t]]$, the adjunction of $\frac{\pi}{\chi_{\vec{m}}}$ in the right hand side of (4) is the base change of a localization corresponding to a toric blow up of $R\times U_{\tau}$, with respect to
(6) $\hat{\tau}^{\prime}:=\tau^{\prime}+\mathbb{R}_{\geq 0}(\vec{0},1).$
In general, we have a natural (birational) map
$b\colon{\overline{\rm ST}}_{R}(\tau^{\prime})\to{\rm Spec}(R)$
and a canonical open embedding ${\rm Spec}(K)\hookrightarrow{\overline{\rm ST}}_{R}(\tau^{\prime}),$ where $K$ denotes the fraction field of the valuation ring $R$. We denote the one dimensional strata of $\widetilde{\rm ST}_{R}(\tau^{\prime})\simeq\mathbb{G}_{m}(\mathbbm{k})$ corresponding to $\tau^{\prime}$ (or its image in ${\overline{\rm ST}}_{R}(\tau^{\prime})$) as $V(\tau^{\prime})$.
@@
Definition 3.3 (Higher $\Theta$-strata cf., [HL14, §2.1] for $r=1$ case).
Consider the mapping stack ${\rm Map}(\Theta_{\tau},\mathcal{M})$ of $\Theta_{\tau}$ to $\mathcal{M}$, as an algebraic stack, after [AHR20, 5.10, 5.11], [AHR19, 6.23], [HLP23, 5.1]. Extending [HL14, Definition 2.1.1], a higher $\Theta$-stratum of type $\tau$ and rank $r$ in $\mathcal{M}$ consists of a union of connected components $\mathcal{Z}^{+}\subset{\rm Map}(\Theta_{\tau},\mathcal{M})$ such that the natural evaluation morphism ${\rm ev}_{(1,\cdots,1)}\colon\mathcal{Z}^{+}\to\mathcal{M}$ is a closed embedding. Note that, if $\tau$ is a regular simplicial one and we fix an isomorphism $\Theta^{r}\simeq\Theta_{\tau}$ for a rational polyhedral cone $\tau$, we can regard it as an open closed substack of ${\rm Filt}^{r}(\mathcal{M}):={\rm Map}(\Theta^{r},\mathcal{M})$.
Furthermore, if $\mathcal{M}$ is quotient stack of the form $[M/G]$ where $M$ is an algebraic $\mathbbm{k}$-scheme $M$ and $G$ is an algebraic $\mathbbm{k}$-group, we often write ${\rm ev}_{(1,\cdots,1)}(\mathcal{Z}^{+})=[Z^{+}/G]$ for some $G$-invariant closed subscheme $Z^{+}\subset M$.
Now, we are ready to introduce the generalization of elementary modification of [AHLH23, §3].
Definition 3.4 (Elementary ($\tau$-)modification cf., [AHLH23, §3.5, Definition 3.36] for $r=1$ case).
An extended elementary ($\tau$-)modification of $f\colon{\rm Spec}(R)\to\mathcal{M}$ is the data of a morphism $\overline{f}\colon\overline{\rm ST}_{R}(\tau)\to\mathcal{M}$ along with an isomorphism $f\simeq\overline{f}|_{\rm Spec(R)}$. Here, to make sense of the restriction symbol, $\rm Spec(R)$ is regarded as a localization of $\overline{{\rm ST}}_{R}(\tau^{\prime})$ since we have $\mathop{\mathrm{Spec}}\nolimits(R)=[\mathop{\mathrm{Spec}}\nolimits R[M]/T]$.
@@
Suppose $(R,v=v_{R})$ is a valuation local ring, such that the associated graded ring $gr_{v}(R):=\oplus_{w\in{\rm Im}(v)}\{a\in R\mid v(a)\geq w\}/\{a\in R\mid v(a)>w\}$ is isomorphic to a semigroup ring over a field $\mathbbm{k}$ and a submonoid of $\mathbb{R}^{r}$ with some $r\in\mathbb{Z}_{>0}$ as a graded ring (which holds e.g. if $R$ is of the form $\mathbbm{k}[[\Gamma_{\geq 0}]]$ for a field $\mathbbm{k}$ in Definition 2.1) and $\xi,N,M_{\xi\geq 0}$ are as Notation 1. Now, we define a ring $R[[M_{\xi\geq 0}]]$ as
$\{\sum_{i=1,2,\cdots}a_{i}t^{m_{i}}\mid a_{i}\in R,m_{i}\in M_{\xi\geq 0},m_{1}<m_{2}<\cdots,m_{i}\to\infty(i\to\infty)\},$
@@
3.2. Higher $\Theta$-semistable reduction theorem
Below, we generalize the semistable reduction type theorem in [AHLH23, §6.1], which was for $r=1$ case. Morally speaking, it says that a one parameter family which strictly degenerates into a higher $\Theta$-strata, there is a canonical modification of the family to avoid a higher $\Theta$-strata i.e., with “less degeneration”. This is a version of semistable reduction type theorem, and later Theorem 3.18 is its iterated version along a certain kind of stratification (higher $\Theta$-stratification cf., Definition 3.17).